In three years, Chad will be three times my present age. I will then be half as old as he. How old am I now?
Answer by usyim88hk(158)

Let x be the age of "I" now
x = my persent age
x+3 = my age after 3 years
3x = Chad's age after 3 years
x+3 = (1/2)3x
(my age after three years is half of Chad's age after three years)
x+3 = 1.5x
3 = 0.5x
x = 6
So now I am 6 years old